# Risk Occupancy: A New and Efficient Paradigm through Vehicle-Road-Cloud Collaboration

### v1

@@ I Introduction
Traffic Scene
4D Risk Occupancy
Figure 1: When the green light comes on and a pedestrian runs through the road, the white ICV fails to recognize the runner due to the large truck. Vehicle-Road-Cloud collaboration can help ICV access the risk occupancy conditions in about 3 seconds, perceiving potential risks in advance.
At present, the intelligent driving industry extensively utilizes 3D object detection[1],[2],[3] or Bird’s Eye View (BEV) object detection[4],[5] in conjunction with local map construction[6],[7],[8],[9]. The intermediate-layer features or result data produced are employed to support subsequent planning algorithms[10],[11],[12],[13]. In contrast to conventional perception paradigms, 4D risk occupancy perception can unify the above two, expressing lane lines and instance targets as corresponding risk occupancy situations and supporting subsequent planning tasks.
This study integrates the conventional driving risk field with the idea of OGM within the Vehicle-Road-Cloud collaborative architecture, offering a novel global perception approach essential for ICV operation in mixed traffic scenarios.
In the field of driving risk field, scholars have proposed the driving safety field model method to describe driving risks, which can be traced back to the work of Gerdes who proposed a driving assistance system based on the APF paradigm [14]. This method uses the field strength to construct an intuitive and unified driving environment, which can describe the inherent constraints of the road environment and collision risks. With the development of related theories, the construction method of the driving risk field model has become more complex, with more factors and methods participating in the risk field construction process. Wang’s team established a driving risk field model composed of driver-vehicle-road factors[15],[16]. Many scholars continue to explore in the field of driving risk field research. For example, Han and others [17] expanded the time dimension on the two-dimensional space, applying a three-dimensional spatiotemporal risk field to trajectory planning. Tian et al.[18] improved the driving risk field for path planning and improved the A* algorithm [19] by adaptively expanding the direction and loss function. In addition, in the research [20], the impact of vehicle motion direction on driving risk and the relationship between potential energy and force were considered, and the driving risk field model was improved, compared, and verified in the following car model. Wang et al.[21] constructed static and dynamic risk matrices for path planning and trajectory planning.
@@
Occupancy grid mapping encapsulates information such as drivable areas and road obstacles, which is an important component of intelligent driving understanding of road scenarios. There have been many related studies over the years. For example, Wang and others[28] used particle filtering and multi-target tracking to achieve dynamic occupancy grid map construction, optimizing the accurate modeling of dynamic environments and target tracking. Zhao and others[29] optimized particle sampling, resampling, and map construction stages, and proposed an offline grid map resolution optimization scheme to improve map resolution while ensuring the real-time nature of the algorithm. Li and others[30] proposed a multi-vehicle collaborative local mapping method based on the fusion of occupancy grid maps, using genetic algorithms to optimize the objective function, and enhancing perception association ability through inter-vehicle relative posture estimation strategies. Toda[31] proposed a path planning algorithm based on multi-resolution occupancy grid maps, effectively reducing the consumption of computing resources in the process of finding feasible paths. The OGM algorithms for vehicle perception also include [32],[33],[34],[35],[36], etc. But the environmental perception methods based on OGM have not made a unified risk assessment of the driving environment.
@@ II-C The Principle of 4D Risk Occupancy
The principle of 4D Risk Occupancy involves accumulating the risk values corresponding to multiple dynamic and static factors within the perceptual range of the sampling points to obtain the final road risk value. This process involves traversing the collection of sampling points and calculating the risk values.
@@ 19             $degree\_risk\leftarrow degree\_risk+risk\_value_{stat}\times weight_{stat}$;
During the assessment process, this paper first defines the perceptual range of the sampling points and adopts corresponding interaction and calculation strategies based on the kind (static or dynamic) of the objects within that range. For dynamic objects, the ETA value needs to be calculated, and the potential collision risk is determined based on the ETA value. For static objects, a fixed risk value and weight are assigned directly according to the type of static object. The pseudo-code for the 4D Risk Occupancy calculation principle is shown in Algorithm 1.
@@ III A Local Path Planning Approach Based on 4D Risk Occupancy
The RRT [37] and RRT* [38] and their variant algorithms belong to the category of sampling-based tree search path methods, suitable for high-dimensional spaces and complex environments, such as [39]. RRT explores the feasible path space through random sampling and tree expansion, capable of handling dynamic obstacles and complex constraints. In terms of combining risk fields, [17] expanded the temporal dimension, constructed a three-dimensional risk field STRF, and developed a trajectory planning algorithm based on it. Wang and others [40] proposed a trajectory planning method for autonomous driving vehicles that combines potential field algorithms and MPC, proposing a human-centered AV control strategy in simulated environments, aiming to balance safety, driving comfort, and smoothness.
@@ IV-B Vehicle-Road-Cloud Collaboration Application Architecture
step=1
step=2
step=3
step=4
@@ V Analysis of Experimental Results
Scheme1
Scheme2
Scheme3

### v2

@@ I Introduction
(a) Traffic Scene
(b) 4D Risk Occupancy
Figure 1: When the green light comes on and a pedestrian runs across the road, the white ICV fails to recognize the runner due to the truck on the right front. Vehicle-Road-Cloud collaboration enables ICV to access the risk occupancy conditions for the next about 3 seconds, perceiving potential risks in advance.
At present, the intelligent driving industry extensively utilizes 3D object detection [1],[2],[3] or Bird’s Eye View (BEV) object detection [4],[5] in conjunction with local map construction [6],[7],[8],[9]. The intermediate-layer features or result data produced are employed to support subsequent planning algorithms [10],[11],[12],[13]. In contrast to conventional perception paradigms, 4D risk occupancy perception paradigm can unify the above two, expressing lane lines and instance targets as corresponding risk occupancy situations and supporting subsequent planning tasks.
This study integrates the conventional driving risk field with the idea of OGM within the Vehicle-Road-Cloud collaborative architecture, offering a novel global perception approach essential for ICV operation in mixed traffic scenarios. In the field of driving risk field, it can be traced back to the work of Gerdes who proposed a driving assistance system based on the APF paradigm [14]. This method uses the field strength to construct an intuitive and unified driving environment, which can describe the inherent constraints of the road environment and collision risks. With the development of related theories, the construction method of the driving risk field model has become more complex, with more factors and methods participating in the risk field construction process. Wang’s team established a driving risk field model composed of driver-vehicle-road factors [15],[16]. Many scholars continue to explore in the field of driving risk field research. For example, Han expanded the time dimension on the two-dimensional space, applying a three-dimensional spatiotemporal risk field to trajectory planning [17]. Tian et al. [18] improved the driving risk field for path planning and improved the A* algorithm [19] by adaptively expanding the direction and loss function. In addition, in the research [20], the impact of vehicle motion direction on driving risk and the relationship between potential energy and force were considered, and the driving risk field model was improved, compared, and verified in the following car model. Wang et al. [21] constructed static and dynamic risk matrices for path planning and trajectory planning.
@@
Occupancy grid mapping encapsulates information such as drivable areas and road obstacles, which is an important component of intelligent driving understanding of road scenarios. There have been many related studies over the years. For example, Wang and others [28] used particle filtering and multi-target tracking to achieve dynamic occupancy grid map construction, optimizing the accurate modeling of dynamic environments and target tracking. Zhao and others [29] optimized particle sampling, resampling, and map construction stages, and proposed an offline grid map resolution optimization scheme to improve map resolution while ensuring the real-time nature of the algorithm. Li [30] proposed a multi-vehicle collaborative local mapping method based on the fusion of occupancy grid maps, using genetic algorithms to optimize the objective function, and enhancing perception association ability through inter-vehicle relative posture estimation strategies. Toda [31] proposed a path planning algorithm based on multi-resolution occupancy grid maps, effectively reducing the consumption of computing resources in the process of finding feasible paths. The OGM algorithms for vehicle perception also include [32],[33],[34],[35],[36], etc. But the environmental perception methods based on OGM have not made a unified risk assessment of the driving environment.
@@ II-C The Principle of 4D Risk Occupancy
The principle of 4D Risk Occupancy involves accumulating the risk values corresponding to multiple dynamic and static factors within the perceptual range of the sampling points to obtain the final road risk value. This process involves traversing the collection of sampling points and calculating the risk values. During the assessment process, this paper first defines the perceptual range of the sampling points and adopts corresponding interaction and calculation strategies based on the kind (static or dynamic) of the objects within that range. For dynamic objects, the ETA value needs to be calculated, and the potential collision risk is determined based on the ETA value. For static objects, a fixed risk value and weight are assigned directly according to the type of static object. The pseudo-code for the 4D Risk Occupancy calculation principle is shown in Algorithm 1.
@@ III A Local Path Planning Approach Based on 4D Risk Occupancy
The RRT [37] and RRT* [38] algorithms, along with their variants, are sampling-based tree search methods suitable for complex, high-dimensional environments [39]. They efficiently explore feasible paths through random sampling and tree expansion, managing dynamic obstacles and constraints. A study [17] has expanded the temporal dimension by constructing a 3D STRF for trajectory planning. Wang and colleagues [40] have proposed a trajectory planning method for autonomous vehicles, combining potential fields with MPC, and developed a human-centered control strategy to balance safety, comfort, and smoothness in simulated environments.
@@ IV-B Vehicle-Road-Cloud Collaboration Application Architecture
Right Refer to caption Refer to caption Refer to caption Refer to caption
Right Refer to caption Refer to caption Refer to caption Refer to caption
Straight Refer to caption Refer to caption Refer to caption Refer to caption
Left Refer to caption Refer to caption Refer to caption Refer to caption
Left Refer to caption step=1 Refer to caption step=2 Refer to caption step=3 Refer to caption step=4
@@ V Analysis of Experimental Results
(a) Scheme1
(b) Scheme2
(c) Scheme3